The line y = kx +b passes through points A (–1, 3) and B (2, –1). Write the equation of the line.
Answer by fractalier(6550)

Couple of ways to do this...let's find the slope first...
k (or m) = (3 - (-1))/(-1 - 2) = -4/3
Now plug into y = kx + b to find b...
3 = (-4/3)(-1) + b
3 = 4/3 + b
b = 5/3
so that
y = (-4/3)x + 5/3